# On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization

### 19 Mar 2026 v1

@@ 1 Introduction
At first glance, this appears plausible since if the change points were known in advance, one could simply restart a stationary satisficing algorithm on each segment, leading to an $\mathcal{O}(L)$ satisficing regret.
@@ 1.1 Our Contributions
  • In Section 4, we establish an $\Omega\bigl(L\log(T/L)\bigr)$ lower bound on the satisficing regret for nonstationary $K$-armed bandits. We then propose Algorithm 1, which achieves an almost matching upper bound of $\mathcal{O}(L\log T)$. Together, these results show that the additional $\log T$ factor is unavoidable in general when the change points are unknown, thereby ruling out the conjectured $\mathcal{O}(L)$ satisficing regret guarantee.
  • To establish the lower bounds above, we extend recent interactive Fano techniques for bandit regret lower bounds (Chen et al. 2024, Shufaro et al. 2024) from stationary to nonstationary environments. More importantly, we introduce a post-interaction reference framework that strictly generalizes the classical Fano approach and is particularly well suited to nonstationary bandits. The key feature is that an auxiliary random variable (the “reference”) may influence the data-generating process and is revealed to the estimator after the interaction. This leads to two complementary principles for lower-bounding regret (Section 3.2): (i) reducing a portion of the regret to an identification error under a suitable separation condition, and (ii) interpreting a portion of the regret as an information cost. We further emphasize that this framework may extend to broader classes of nonstationary bandit problems and may be of independent interest.
  • In Section 5, we complement the general negative result by identifying a special case in which constant satisficing regret is still achievable. Specifically, when there exists an always-realizable arm and no satisficing arm ever becomes non-satisficing, the stationary algorithms of Michel et al. (2022) can be applied directly in the nonstationary setting, yielding constant satisficing regret that is independent of both the horizon $T$ and the number of stationary segments $L$.
@@ 1.2 Related Work
Nonstationary bandits.
Most work on nonstationary bandits evaluates performance through dynamic regret, which compares the learner with the best arm at each round. For piecewise-stationary environments with at most $L$ segments, sliding-window and change-point-based methods achieve $\widetilde{\mathcal{O}}(\sqrt{LT})$ dynamic regret (Garivier and Moulines 2011, Auer et al. 2019). Under a variation-budget model with total drift $D$, the optimal rate is $\widetilde{\mathcal{O}}(D^{1/3}T^{2/3})$ (Besbes et al. 2014, Cheung et al. 2019), and Wei and Luo (2021) obtain the unified guarantee $\widetilde{\mathcal{O}}(\min\{\sqrt{LT},D^{1/3}T^{2/3}\})$. These results address a stronger benchmark than ours: dynamic regret competes with the instantaneous best arm, whereas our objective only requires selecting an arm whose mean exceeds a fixed threshold.
Stationary satisficing regret.
Our work is most closely related to the literature on threshold-based satisficing regret in stationary bandits. Michel et al. (2022), Rouyer et al. (2024), and Feng et al. (2024) study cumulative satisficing regret under a fixed threshold and show that, under realizability, constant regret is achievable in stationary settings. Our paper shows that this phenomenon changes sharply under nonstationarity: once the environment may change at unknown times, constant satisficing regret is no longer achievable in general.
Threshold-based pure exploration and other satisficing notions.
In stationary settings, threshold-based objectives are also studied in pure exploration under the name thresholding bandits (Locatelli et al. 2016, Mukherjee et al. 2017). These works aim to identify the arms whose means exceed a threshold and are evaluated by probability of error or sample complexity, rather than cumulative satisficing regret. Although such methods can identify a satisficing arm with high probability, a naive explore–then–exploit reduction is not sufficient to obtain optimal cumulative satisficing regret even in stationary problems (Michel et al. 2022, Feng et al. 2024).
@@
For nonstationary bandits, as highlighted in Liu et al. (2023b, a), the standard notion of dynamic regret can steer algorithm design toward overly aggressive exploration; it can therefore be more meaningful to compete against a satisficing benchmark. Min and Russo (2023) study nonstationary bandits in a Bayesian model and introduce a notion of satisficing based on competing with a satisficing action sequence $A^{\dagger}$ (Min and Russo 2023, Sec. 7). Their ‘satisficing regret” is defined as regret relative to $A^{\dagger}$. In contrast, we study a threshold-based notion of satisficing regret, defined as the cumulative shortfall from a fixed threshold $S$. Our formulation does not require that any arm achieves mean exactly $S$ at each time. Even if one sets $A^{\dagger}$ to be a realizable arm sequence, the two regret notions are not directly comparable: regret relative to $A^{\dagger}$ can be negative at times when the learner selects an arm with strictly larger mean than $A_{t}^{\dagger}$, whereas our satisficing regret is nonnegative by definition. Finally, Min and Russo (2023) establishes asymptotic lower bounds for the classical dynamic regret, while we provide minimax-optimal finite-time guarantees for threshold-based satisficing regret.
@@ 1.3 Notation
For integers $M\geq 1$, define $M]:=\{1,2,\ldots,M\}$. For integers $0\leq N\leq M$, define $N,M]:=\{N,N+1,\ldots,M\}$. For a buffer $\mathcal{B}=(b_{1},\ldots,b_{m})$ of length $m$, we write $\mathcal{B}[s]:=b_{s}$ for its $s$-th entry, $s\in[m]$. For a set $\mathcal{A}$, we write $|\mathcal{A}|$ for its cardinality, and $\mathrm{Unif}(\mathcal{A})$ for the uniform distribution on $\mathcal{A}$. For random variables $(X,Y)$ with joint law $P_{X,Y}$, the mutual information is
@@ 2 Problem Formulation
Let $L:=L[1,T]$ denote the total number of stationary segments, and let
$1=T_{0}<T_{1}<\cdots<T_{L}=T+1$
be the associated change-points that partition the horizon into $L$ stationary segments, where segment $\ell\in\{0,\dots,L-1\}$ is the interval $T_{\ell},\,T_{\ell+1}-1]$. A (possibly randomized) policy $\pi$ selects an action $a_{t}\in[K]$ at time $t$, adapted to the natural filtration $\mathcal{F}_{t-1}:=\sigma(a_{1},r_{1},\dots,a_{t-1},r_{t-1})$. Let $\Pi$ denote the class of all such policies. Conditional on $a_{t}$, the learner observes a reward $r_{t}=\mu_{t}(a_{t})+\xi_{t},$ where $(\xi_{t})_{t\geq 1}$ are independent, zero-mean, $1$-sub-Gaussian noises.
Satisficing Regret.
@@
where $\nu$ denotes the underlying nonstationary environment and $\mathbb{E}_{\pi,\nu}$ is over the policy’s randomness and the observation noise.
Satisficing nonstationary bandits.
For each segment $\ell\in\{0,\dots,L-1\}$, let $\mu_{a,\ell}:=\mu_{T_{\ell}}(a)$ denote the mean reward of arm $a\in[K]$ on segment $\ell$, and define the segment-wise optimal mean $\mu_{\ell}^{*}:=\max_{a\in[K]}\mu_{a,\ell}$. Throughout, we impose the following realizability condition.
@@
Assumption 1 is standard in studies of satisficing regret in the stationary setting (Michel et al. 2022, Feng et al. 2024). We refer to the nonstationary bandit problem under Assumption 1, together with the objective (2), as the nonstationary satisficing bandit problem.
@@ 3 A Fano-Based Framework for Nonstationary Bandit Problems
Fano’s lemma (Reza 1994, Theorem 2.10) is a classical tool for minimax lower bounds in multi-hypothesis testing. Let $\mathcal{V}$ be a finite index set, let $\nu\sim\mathrm{Unif}(\mathcal{V})$ be a latent instance index, and let $Z$ denote the data generated by an algorithm interacting with the instance indexed by $\nu$. Fano’s lemma lower bounds the best achievable identification error in terms of the mutual information $I(\nu;Z)$:
@@
Recently, building on (3), Chen et al. (2024), Shufaro et al. (2024) developed general Fano-type frameworks for proving regret lower bounds in stationary bandit problems. In Section 3.1, we review how these works leverage (3) to derive interactive Fano methods for stationary bandits.
Conditional Fano.
It is often useful to apply Fano’s lemma in a structured setting where the index takes the form $\nu=(u,b)\in\mathcal{U}\times\mathcal{B}$. Here the “reference” variable $b$ is independent of $u$, the observation $Z$ may depend on $b$, and $b$ is revealed to the estimator before it outputs $\widehat{u}$. A conditional form of Fano’s inequality (Scarlett and Cevher 2019, Theorem 3) then follows directly from (3):
@@
In Section 3.2, building on (4), we develop a post-interaction reference framework which is particularly convenient for nonstationary bandits and yields two complementary principles for deriving regret lower bounds.
@@ 3.1 Classical Fano Reduction: Small Regret Implies Correct Identification
Draw an environment index $\nu\sim\mathrm{Unif}(\mathcal{V})$. Fix any policy $\pi$, and let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ denote the full interaction transcript obtained by running $\pi$ for $T$ rounds in environment $\nu$. Let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$, and let $\bar{P}:=\frac{1}{|\mathcal{V}|}\sum_{\tau\in\mathcal{V}}P_{\tau}$ be the resulting mixture distribution. By the definition of mutual information and the convexity of $D_{\mathrm{KL}}(P\|\cdot)$ in its second argument (Cover 1999), we have
@@
In stationary bandits, a standard route to regret lower bounds is to reduce small regret to accurate identification of the underlying instance. Concretely, one constructs an instance family $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ satisfying a separation condition at level $\Delta$ (see, e.g., Shufaro et al. 2024, Prop. 3.2 or the more general formulation in Chen et al. 2024, Eq. (10)), which enables the construction of an estimator $\widehat{\nu}(Z)$ from the transcript $Z$ such that
@@
where $\mathrm{Reg}_{T}$ denotes the regret of interest. By Markov’s inequality and (6), the Bayesian regret under the uniform prior satisfies
@@ 3.2 Post-Interaction Reference and Two Complementary Approaches
In nonstationary settings, the reduction (7) typically breaks down: unlike the stationary case, there is no canonical separation condition that directly links regret to identifying the latent instance, since the underlying structure may vary over time. Motivated by the conditional bound (4), we introduce an interaction protocol with a post-interaction reference.
Protocol.
Draw an environment index $(u,b)\sim\mathrm{Unif}(\mathcal{V})$, where $\mathcal{V}:=\mathcal{U}\times\mathcal{B}$. Under the prior $\mathrm{Unif}(\mathcal{V})$, the random variables $u$ and $b$ are independent. Under the (unknown) environment $(u,b)$, let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ denote the interaction transcript, and let $P_{u,b}$ denote the law of $Z$ conditional on $(u,b)$. The key feature is that $b$ (the reference) is revealed to the estimator only after $Z$ has been generated. When $b$ is uninformative, this reduces to the classical formulation in Section 3.1.
Conditional versions of Fano’s lemma have been widely studied in statistics and information theory (Atia and Saligrama 2012, Scarlett and Cevher 2016, 2019), but have not, to our knowledge, been systematically developed as a general lower-bound mechanism for nonstationary bandits. We show that a post-interaction reference enables two complementary approaches for proving regret lower bounds.
Approach I: coordinate-wise identification error.
Analogous to (5), conditioning on $b$ yields
@@
A particularly useful instantiation is when $\mathcal{V}$ has a product structure $\mathcal{V}=\mathcal{V}_{1}\times\cdots\times\mathcal{V}_{m}$. For any coordinate $i\in[m]$, we take
@@
This yields a coordinate-wise identification bound which can be paired with a lower bound on a partial regret component:
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\lesssim\mathbb{E}\big[\mathrm{Reg}_{T}^{(i)}(\pi,\nu)\big].$ (10)
Throughout Section 3, the notation $\lesssim$ may hide constants that depend on $T$ and $\{P_{\nu}\}_{\nu\in\mathcal{V}}$. An intuitive interpretation of (8) and (10) is as follows: under an appropriate separation condition for a carefully constructed family of nonstationary instances, once the reference is revealed, any policy $\pi$ must incur a nontrivial coordinate-wise identification error. This nontrivial lower bound on the identification error can then be translated into a lower bound on the corresponding partial regret. If, moreover, the total regret decomposes as $\mathbb{E}[\mathrm{Reg}_{T}(\pi;\nu)]=\sum_{i=1}^{m}\mathbb{E}[\mathrm{Reg}_{T}^{(i)}(\pi;\nu)]$, then summing coordinate-wise bounds and optimizing over $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ yields Bayesian regret lower bounds. Since (10) is more general than (7)—both in allowing a broader separation condition and in relating identification to partial regret through the post-interaction reference—Approach I can be viewed as a strict extension of the interactive Fano methods for stationary bandits outlined in Section 3.1.
This coordinate-wise viewpoint is closely related to Assouad’s method for (passive) statistical estimation (Yu 1997). In the bandit setting, Chen et al. (2024) develop a general framework for stationary problems that recovers Assouad-type arguments. We show that the reference-aided coordinate-wise strategy extends naturally to nonstationary bandits: it yields the optimal dependence on the number of stationary segments $L$, and we leverage it to prove Theorem 1.
Approach II: coordinate-wise information cost .
We again adopt the product structure (9). Conditional on $(\mathcal{F}_{t-1},b)$, the action $a_{t}$ is drawn using the policy’s internal randomness, which is independent of $u$. Hence $I(u;a_{t}\mid\mathcal{F}_{t-1},b)=0$, and the chain rule yields
@@
The strength of (11) is that it yields regret lower bounds without requiring a direct implication of the form “(under an appropriate separation condition) small (partial) regret $\Rightarrow$ correct (coordinate-wise) identification,” an implication that is used both in the prior interactive Fano methods reviewed in Section 3.1 and in our Approach I, but that can fail in certain nonstationary constructions. Instead, we adopt an information cost perspective: we construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ and an estimator $\widehat{u}(Z,b)$ such that for some $\rho\in(0,1)$,
@@
An intuitive interpretation of (11) and (12) is as follows: under any policy $\pi$, once the reference is revealed, the coordinate-wise identification error admits a nontrivial upper bound. Then, by (11), the learner must have acquired a sufficient amount of information about the underlying coordinate; in our construction, acquiring this information necessarily incurs some partial regret. In many nonstationary constructions, once $b$ is revealed, information about $u$ is acquired only through actions that incur regret, which makes (12) easier to verify than a direct reduction (10). We use this approach to prove Theorem 2.
We emphasize that, although classical KL change-of-measure arguments also link regret to information acquisition, they differ in an essential way from our approach. In the stationary setting, such arguments typically compare the laws of the interaction transcript under a small family of alternative environments and convert the resulting KL divergences into regret lower bounds (Bubeck et al. 2012, Garivier et al. 2019). Related ideas also appear in nonstationary bandits (Garivier and Moulines 2011). However, these arguments are usually organized around pairwise comparisons and are expressed directly in terms of KL divergence. By contrast, our lower-bound construction is genuinely multi-hypothesis, and it has a product structure across coordinates and is naturally formulated in terms of conditional mutual information. For this reason, a direct pairwise change-of-measure argument does not by itself provide the coordinate-wise decomposition and information-cost analysis required in our setting. We discuss this distinction further in the next section, where we explain why such classical arguments do not directly yield the desired logarithmic lower bounds for nonstationary satisficing bandits.

### 25 Apr 2026 v2

@@ 1 Introduction
At first glance, this seems plausible: if the change points were known in advance, one could simply restart a stationary satisficing algorithm on each segment, leading to an $\mathcal{O}(L)$ satisficing regret.
@@ 1.1 Our Contributions
  • In Section 4, we establish an $\Omega\bigl(L\log(T/L)\bigr)$ lower bound on the satisficing regret for nonstationary $K$-armed bandits. We then propose Algorithm 1, which achieves an almost matching upper bound of $\mathcal{O}(L\log T)$. Together, these results show that the additional $T$-dependent $\log T$ factor is unavoidable in general when the change points are unknown, thereby ruling out the conjectured $\mathcal{O}(L)$ satisficing regret guarantee.
  • To establish the lower bounds above, we extend recent interactive Fano techniques for bandit regret lower bounds (Chen et al. 2024, Shufaro et al. 2024) from stationary to nonstationary environments. More importantly, we introduce a post-interaction reference framework that strictly generalizes the classical Fano approach to make it well suited to nonstationary bandits. In this framework, a key distinction compared to the Fano techniques is that part of the (initially) unknown environment parameters (the “reference”) are revealed to the estimator after the entire horizon of length $T$. This leads to two novel complementary paths in lower bounding the regret (Section 3.2): (i) reducing a portion of the regret to an identification error under a suitable separation condition (used in establishing the lower bound when $L\geq 3$); and (ii) interpreting a portion of the regret as an information cost (used for $L=2$). We also note that this framework may extend to broader classes of nonstationary bandit problems and may be of independent interest.
  • In Section 5, we complement the general negative results by identifying a special case in which constant satisficing regret is achievable. Specifically, when there exists an always-realizable arm and no satisficing arm ever becomes non-satisficing, following the results in Michel et al. (2022) can yield a constant satisficing regret that is independent of both $T$ and $L$.
@@ 1.2 Related Work
Nonstationary Bandits. Most work on nonstationary bandits evaluates performance through dynamic regret, which compares the learner with the best arm at each round. For piecewise-stationary environments with at most $L$ segments, sliding-window and change-point-based methods achieve $\widetilde{\mathcal{O}}(\sqrt{LT})$ dynamic regret (Garivier and Moulines 2011, Auer et al. 2019). Under a variation-budget model with total drift $D$, the optimal rate is $\widetilde{\mathcal{O}}(D^{1/3}T^{2/3})$ (Besbes et al. 2014, Cheung et al. 2019), and Wei and Luo (2021) obtain the unified guarantee $\widetilde{\mathcal{O}}(\min\{\sqrt{LT},D^{1/3}T^{2/3}\})$. These results address a stronger benchmark than ours: dynamic regret competes with the instantaneous best arm, whereas our objective only requires selecting an arm whose mean exceeds a fixed threshold.
Stationary Satisficing Regret. Our work is most closely related to the literature on threshold-based satisficing regret in stationary bandits. Michel et al. (2022), Rouyer et al. (2024), and Feng et al. (2024) study cumulative satisficing regret under a fixed threshold and show that, under realizability, constant regret is achievable in stationary settings. Our paper shows that this phenomenon changes sharply under nonstationarity: once the environment may change at unknown times, constant satisficing regret is no longer achievable in general.
Other Satisficing Notions. In stationary settings, threshold-based objectives have also been studied in pure exploration under the name thresholding bandits (Locatelli et al. 2016, Mukherjee et al. 2017). There, the goal is to identify arms above a threshold, with performance measured by error probability or sample complexity rather than cumulative satisficing regret. Moreover, a naive explore–then–exploit reduction does not yield optimal cumulative satisficing regret even in the stationary case (Michel et al. 2022, Feng et al. 2024).
@@
For nonstationary bandits, as highlighted in Liu et al. (2023b, a), the standard notion of dynamic regret can steer algorithm design toward overly aggressive exploration; it can therefore be more meaningful to compete against a satisficing benchmark. Min and Russo (2023) study nonstationary bandits in a Bayesian model and introduce a notion of satisficing based on competing with a satisficing action sequence $A^{\dagger}$ (Min and Russo 2023, Sec. 7). Their “satisficing regret” is defined as regret relative to $A^{\dagger}$. In contrast, we study a threshold-based notion of satisficing regret, defined as the cumulative shortfall from a fixed threshold $S$. Our formulation does not require that any arm achieves mean exactly $S$ at each time. Even if one sets $A^{\dagger}$ to be a realizable arm sequence, the two regret notions are not directly comparable: regret relative to $A^{\dagger}$ can be negative at times when the learner selects an arm with strictly larger mean than $A_{t}^{\dagger}$, whereas our satisficing regret is nonnegative by definition. Finally, Min and Russo (2023) establishes asymptotic lower bounds for the classical dynamic regret, while we provide minimax-optimal finite-time guarantees for threshold-based satisficing regret.
@@ 1.3 Notation
For integers $M\geq 1$, define $M]:=\{1,2,\ldots,M\}$. For integers $0\leq N\leq M$, define $N,M]:=\{N,N+1,\ldots,M\}$. For a buffer $\mathcal{B}=(b_{1},\ldots,b_{m})$ of length $m$, we write $\mathcal{B}[s]:=b_{s}$ for its $s$-th entry, $s\in[m]$. For a set $\mathcal{A}$, we use $|\mathcal{A}|$ to denote its cardinality, and $\mathrm{Unif}(\mathcal{A})$ for the uniform distribution on $\mathcal{A}$. For random variables $(X,Y)$ with joint law $P_{X,Y}$, the mutual information is
@@ 2 Problem Formulation
Let $L:=L[1,T]$ denote the total number of stationary segments and $1=T_{0}<T_{1}<\cdots<T_{L}=T+1$ be the associated change-points that partition the horizon into $L$ stationary segments, where segment $\ell\in\{0,\dots,L-1\}$ is the interval $T_{\ell},\,T_{\ell+1}-1]$. A (possibly randomized) policy $\pi$ selects an action $a_{t}\in[K]$ at time $t$, adapted to the natural filtration $\mathcal{F}_{t-1}:=\sigma(a_{1},r_{1},\dots,a_{t-1},r_{t-1})$. Let $\Pi$ denote the class of all such policies. Conditional on $a_{t}$, the learner observes a reward $r_{t}=\mu_{t}(a_{t})+\xi_{t},$ where $(\xi_{t})_{t\geq 1}$ are independent, zero-mean, $1$-sub-Gaussian noises. For each segment $\ell\in\{0,\dots,L-1\}$, let $\mu_{a,\ell}:=\mu_{T_{\ell}}(a)$ denote the mean reward of arm $a\in[K]$ on segment $\ell$, and define the segment-wise optimal mean $\mu_{\ell}^{*}:=\max_{a\in[K]}\mu_{a,\ell}$.
@@
where $\nu$ denotes the underlying nonstationary environment and $\mathbb{E}_{\pi,\nu}$ is over the policy’s randomness and the observation noise. Throughout, we focus on the standard realizable case as specified below (Michel et al. 2022, Feng et al. 2024).
@@ 3 A Fano-Based Framework for Nonstationary Bandit Problems
In this section, we detail our new technical development based on Fano’s lemma (Reza 1994, Theorem 2.10) . We start by reviewing its usage for lower bounds in stationary bandits (Section 3.1), and then derive the post-interaction reference framework (Section 3.2). This yields two powerful approaches that are particularly suitable for deriving regret lower bounds for nonstationary bandits.
Fano’s lemma is a classical tool for establishing minimax lower bounds in multi-hypothesis testing. Let $\mathcal{V}$ be a finite index set, $\nu\sim\mathrm{Unif}(\mathcal{V})$ be a latent instance index, and $Z$ be the data generated by an algorithm interacting with the instance indexed by $\nu$, Fano’s lemma lower bounds the probability of being unable to correctly identify $\nu$ with $Z$ in terms of the mutual information $I(\nu;Z)$:
@@
Being initially designed for multi-hypothesis testing, recent works (Chen et al. 2024, Shufaro et al. 2024) build on it to develop a general framework for proving regret lower bounds in stationary bandit problems.
Conditional Fano. Aside from the above, it is often useful to apply Fano’s lemma in a structured setting where the index takes the form $\nu=(u,b)\in\mathcal{U}\times\mathcal{B}$. Here, $b$ is a “reference” variable independent of $u$. The observation $Z$ may depend on $b$, and $b$ is revealed to the estimator before it outputs $\widehat{u}$. A conditional form of Fano’s inequality (Scarlett and Cevher 2019, Theorem 3) then follows directly from (3):
@@ 3.1 Classical Fano Reduction: Small Regret Implies Correct Identification
For a bandit problem, we write out $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ as the full interaction transcript obtained by running $\pi$ for $T$ rounds under the environment indexed by $\nu$. Let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$ and $\bar{P}:=\frac{1}{|\mathcal{V}|}(\sum_{\tau\in\mathcal{V}}P_{\tau})$ be the resulting mixture distribution. By definition of mutual information and the convexity of $D_{\mathrm{KL}}(P\|\cdot)$ in its second argument (Cover 1999), we have
@@
In stationary bandits, a standard route to regret lower bounds is to reduce small regret to accurate identification of the underlying instance $\nu$. Specifically, one constructs an instance family $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ satisfying a separation condition (among different arms’ mean rewards) at level $\Delta$ (see, e.g., Chen et al. 2024, Eq. (10)), which enables the construction of an estimator $\widehat{\nu}(Z)$ from the transcript $Z$ such that
@@
where $\mathrm{Reg}_{T}$ denotes the regret of interest. By Markov’s inequality and (6), the regret under the uniform prior satisfies
@@ 3.2 Post-Interaction Reference and Two Complementary Approaches
In nonstationary bandits, however, the reduction (7) typically fails because the canonical separation condition may not hold due to changes in the arms’ mean rewards. To overcome this, we introduce the post-interaction reference based on the conditional bound (4). To the best of our knowledge, our framework is the first systematic application of the conditional Fano’s method to establish lower bounds in (nonstationary) bandits.
Post-Interaction Reference Protocol. Draw an environment index $(u,b)\sim\mathrm{Unif}(\mathcal{V})$, where $\mathcal{V}:=\mathcal{U}\times\mathcal{B}$. Under the prior $\mathrm{Unif}(\mathcal{V})$, the random variables $u$ and $b$ are independent. Under the (unknown) environment $(u,b)$, we let $P_{u,b}$ denote the law of $Z$ conditional on $(u,b)$. The key feature of the post-interaction reference framework is that $b$ (the reference) is revealed to the estimator only after $Z$ has been generated. This enables the below two complementary approaches for proving regret lower bounds.
Approach I: Coordinate-Wise Identification Error. Analogous to (5), conditioning on $b$ yields
@@
A particularly useful instantiation is when $\mathcal{V}$ has a product structure $\mathcal{V}=\mathcal{V}_{1}\times\cdots\times\mathcal{V}_{m}$. For any coordinate $i\in[m]$, we can take
@@
This coordinate-wise construction is particularly well suited to nonstationary bandits: each factor $\mathcal{V}_{i}$ can be viewed, roughly, as an instance family associated with one stationary segment. This structure is key to obtaining the optimal dependence on the number of stationary segments. In particular, it yields a coordinate-wise identification bound that can be paired with a lower bound on a corresponding partial regret term:
$\inf_{\widehat{u}(Z,b)}\mathbb{P}\bigl(\widehat{u}(Z,b)\neq u\bigr)\lesssim\mathbb{E}\big[\mathrm{Reg}_{T}^{(i)}(\pi,\nu)\big],$ (10)
and the total regret decomposes as $\mathbb{E}[\mathrm{Reg}_{T}(\pi;\nu)]=\sum_{i=1}^{m}\mathbb{E}[\mathrm{Reg}_{T}^{(i)}(\pi;\nu)]$. Throughout this section, the notation $\lesssim$ hides constants that may depend on $T$ and $\{P_{\nu}\}_{\nu\in\mathcal{V}}$.
The intuition for (8) and (10) is as follows: for instances that under an appropriate separation condition on $\mathcal{U}$, once the reference $b\in\mathcal{B}$ is revealed, any policy $\pi$ must incur a nontrivial coordinate-wise identification error. This lower bound on the identification error can in turn be converted into a lower bound on the associated partial regret in (10). Summing them up and optimizing over $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ then yields regret lower bounds. Approach I strictly extends the interactive Fano methods for stationary bandits outlined in Section 3.1: it allows a broader separation condition and links identification to partial regret via the post-interaction reference. It underlies our proof of Theorem 1, which establishes the logarithmic lower bound for $L\geq 3$.
However, as discussed after Theorem 1, when $L=2$ it is unclear how to apply Approach I. To overcome this, we develop a complementary method, again based on the conditional Fano inequality (4) and the post-interaction reference.
Approach II: Coordinate-Wise Information Cost. We again work under the product structure of $\mathcal{V}$, with $\mathcal{U}$ and $\mathcal{B}$ chosen as in (9) for each $i\in[m]$. Conditional on $(\mathcal{F}_{t-1},b)$, the action $a_{t}$ is drawn from the policy’s internal randomness and is therefore independent of $u$. Hence $I(u;a_{t}\mid\mathcal{F}_{t-1},b)=0$, and the chain rule gives
@@
The advantage of (11) is that it yields regret lower bounds without requiring a direct reduction from small (partial) regret to correct (coordinate-wise) identification, a step needed in Approach I but one that may fail in some nonstationary constructions. Instead, we take an information-cost view: we construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$ and an estimator $\widehat{u}(Z,b)$ such that for some $\rho\in(0,1)$,
@@
An intuitive interpretation of (11) and (12) is as follows: once the reference is revealed, any policy $\pi$ must achieve a nontrivial level of coordinate-wise identification accuracy. Thus, (11) yields a lower bound on the information that must be acquired. In our construction, acquiring this information necessarily incurs partial regret. We use this approach to prove Theorem 2, which establishes the logarithmic lower bound for $L=2$ and complements Theorem 1.
Remark 1 (Comparison with the Classical KL-Based Approach).
Classical KL-based arguments also relate regret to information acquisition, but they are inherently pairwise: they compare transcript laws under two alternative environments and convert the resulting KL divergence into a regret lower bound (Bubeck et al. 2012, Garivier et al. 2019, Garivier and Moulines 2011). By contrast, our construction is genuinely multi-hypothesis, via Fano’s inequality, and is therefore better suited to capture the optimal dependence on the number of hypotheses and hence sharper regret lower bounds. Moreover, the relevant information measure in our framework is (conditional) mutual information, which is central to our analysis and can sometimes be linked to regret more directly than KL divergence. We provide a more detailed discussion of this after Theorem 2.